Why dielectric constant effect electrostatic force, what is its significance?
1
Expert's answer
2020-06-25T09:36:12-0400

When dielectric materials is placed between two charges, it polarizes the electric field between them ,thus effects the net electric field and since Columb's force is proportional to electric field ,hence it effects the force, moreover mathematically force between the charges is described by

"F=\frac{1}{4\pi\epsilon_o\epsilon_r}\frac{q_1q_2}{r^2}"

Where,"\epsilon_r" is the dielectric constant of the dielectric medium.

Clearly,

"F\propto\frac{1}{\epsilon_r}"

Hence, the fact.